In triangle ABC, angle B is five times as large as angle C. Angle A
measures 5 degrees more than angle C. Calculate the angles of the triangle(angles A,B,C_. The sum of all angles equal 180 degrees.

Answer by mananth(16946)

let angle c be x
angle b = 5x
angle a = x+5
...
x+x+5+5x= 180
7x+5=180
7x=180-5
7x=175
/7
x=25 deg. ....angle c
angle b = 5x= 125
angle a = x+5 = 30
...
m.ananth@hotmail.ca